# Robust Anomaly Detection via Tensor Chidori Pseudoskeleton Decomposition

### v2

@@ Abstract
Anomaly detection plays a critical role in modern data-driven applications, from identifying fraudulent transactions and safeguarding network infrastructure to monitoring sensor systems for irregular patterns. Traditional approaches—such as distance-, density-, or cluster-based methods, face significant challenges when applied to high-dimensional tensor data, where complex interdependencies across dimensions amplify noise and computational complexity. To address these limitations, this paper leverages Tensor Chidori pseudoskeleton decomposition within a tensor-robust principal component analysis framework to extract low-Tucker-rank structure while isolating sparse anomalies, ensuring robustness to anomaly detection. We establish theoretical analysis of convergence, and estimation error, demonstrating the stability and accuracy of the proposed approach. Numerical experiments on real-world spatiotemporal data from New York City taxi trip records validate the superiority of the proposed method in detecting anomalous urban events compared to existing benchmark methods. The results underscore the potential of Tensor Chidori pseudoskeleton decomposition to enhance anomaly detection for large-scale, high-dimensional data.
@@ I Introduction
To address these limitations, Tensor Chidori pseudoskeleton decomposition offers an alternative approach by selecting representative parts of the data, thereby preserving essential features while reducing redundancy. Tucker Chidori pseudoskeleton decomposition provides a structured decomposition that is both computationally efficient and robust [9].
In this paper, we propose an Robust Principle Component Anlysis algorithm motivated by Tucker Chidori pseudoskeleton decomposition framework [9] tailored for anomaly detection in high-dimensional datasets. By incorporating sparsity and regularization constraints, our method reduces sensitivity to anomalies, enabling more accurate and resilient detection of unusual patterns. The Tucker Chidori pseudoskeleton decomposition framework combines the strengths of Tucker decomposition’s structural insight with pseudoskeleton’s selective feature extraction while enhancing robustness against outliers [9].
@@ I-A Notations and definitions
In the realm of matrix algebra, the pseudoskeleton decomposition technique is a good alternative to SVD [11]. Specifically, this method entails selecting specific columns ${C}$ and rows ${R}$ from a matrix ${X}\in\mathbb{R}^{d_{1}\times d_{2}}$, and constructing a core matrix ${U}={X}(I,J)$. The matrix ${X}$ is then reconstructed through the product ${C}{U}^{\dagger}{R}$, under the condition that $\operatorname{rank}({U})=\operatorname{rank}({X})$. Expanding from matrices to tensors, the initial adaptations of pseudoskeleton decompositions applied a single-mode unfolding to 3-mode tensors [12]. Recent advances have further refined these techniques, introducing the terms Chidori pseudoskeleton decompositions [9].
Theorem 1 [9, Theorem 4.2]).
@@ II Methodology
$\begin{split}\min_{\mathcal{L},\mathcal{S}}&\quad\|\mathcal{T}-\mathcal{L}-\mathcal{S}\|_{\mathrm{F}}\\ \text{subject to}&\quad\mathcal{L}\text{ is low-Tucker-rank and }\mathcal{S}\text{ is sparse}.\end{split}$
The Frobenius norm $\|\mathcal{T}-\mathcal{L}-\mathcal{S}\|_{\mathrm{F}}$ represents the reconstruction error, ensuring that the sum of $\mathcal{L}$ and $\mathcal{S}$ closely approximates $\mathcal{T}$. The constraints on $\mathcal{L}$ and $\mathcal{S}$ are key to achieving meaningful decomposition. The low-Tucker-rank constraint on $\mathcal{L}$ generalizes the concept of low-rank structure to tensors, where the rank is defined across multiple modes rather than a single dimension. This constraint ensures that $\mathcal{L}$ captures only the dominant correlations across spatial and temporal dimensions, filtering out noise while preserving the primary structure. On the other hand, the sparsity constraint on $\mathcal{S}$ ensures that anomalies are localized, representing only a small fraction of the entries in $\mathcal{T}$. This assumption aligns with real-world spatial-temporal datasets, where anomalies like traffic accidents, unusual weather conditions, or security breaches are sparse and isolated.
II-A TRPCA via Tensor Chidori Pseudoskeleton Decomposition
Algorithm 1 TRPCA via Tensor Chidori Pseudoskeleton Decomposition
1:Input: $\mathcal{T}=\mathcal{L}^{\star}+\mathcal{S}^{\star}\in\mathbb{R}^{d_{1}\times\cdots\times d_{n}}$: observed tensor; $(r_{1},\cdots,r_{n})$: underlying Tucker rank of $\mathcal{L}^{\star}$; $\varepsilon$: targeted precision; $\zeta^{(0)},\gamma$: thresholding parameters; $\{|I_{i}|\}_{i=1}^{n},\{|J_{i}|\}_{i=1}^{n}$: cardinalities for sample indices.
@@ II-A1 Step (I): Update Sparse Component $\mathcal{S}$
In this step, we focus on updating the sparse component $\mathcal{S}$, which captures the outliers within the data. To achieve this, we employ a straightforward yet powerful approach: the hard thresholding operator, denoted by $\mathrm{HT}_{\zeta}$. This operator is particularly effective for isolating outlier elements by setting small-magnitude entries to zero, thus retaining only values that exceed a specified threshold, $\zeta$.
The hard thresholding operator $\mathrm{HT}_{\zeta}$ is defined as follows:
@@
This operator $\mathrm{HT}_{\zeta}$ effectively filters out entries with magnitudes less than or equal to $\zeta$, treating them as negligible. By applying this to the tensor $\mathcal{T}$, only values deemed significant (i.e., values exceeding the threshold) remain in the updated sparse component $\mathcal{S}$. Consequently, this approach emphasizes substantial elements while removing noise, thereby enhancing the sparsity of $\mathcal{S}$.
@@ II-A2 Step (II): Update Low-Tucker-rank Component $\mathcal{L}$
In this step, we aim to update the low-Tucker-rank component $\mathcal{L}$, which models the structured, low-rank part of the data tensor. The update process is divided into two key stages: subspace identification and projective reconstruction. To approximate the low-rank structure along each mode, we begin by extracting the mode-$i$ fibers from the residual tensor $\mathcal{T}-\mathcal{S}^{(k)}$, which represents the current estimate of the sparse component subtracted from the observed data tensor. The fibers are assembled into the matrix representation:
@@
This projection aligns the updated factor matrices along mode-$i$ with the estimated low-dimensional subspace, ensuring that the Tucker core captures the dominant variation along this mode. By leveraging QR decomposition and projecting onto the selected subspace, the computational complexity for each mode is reduced from the cubic cost $\mathcal{O}(d_{i}^{3})$ to the more efficient: $\mathcal{O}(d_{i}r_{i}^{2}+r_{i}^{3}),$ where $d_{i}$ is the dimension along mode-$i$, and $r_{i}$ is the target Tucker rank. This reduction is particularly beneficial when the Tucker rank $r_{i}$ is significantly smaller than the mode dimension $d_{i}.$
@@ IV Numerical Experiments
In our experiments, we aggregate the data by counting the number of arrivals per zone over hourly intervals. To ensure statistical significance, we restrict our analysis to 81 central zones, which represent high-traffic areas and exclude zones with minimal activity. This selection reduces noise from sparsely populated zones and provides a more robust representation of NYC’s high-demand regions. With these parameters, we constructed a four-dimensional tensor $\mathbf{Y}$ with dimensions $24\times 7\times 53\times 81$. The modes of this tensor are defined as follows: the first mode corresponds to the 24 hours of a day; the second mode represents the 7 days of the week; the third mode encompasses the 53 weeks of the year; the fourth mode covers the 81 selected central zones in New York City. Thus, each entry in the tensor represents the count of taxi arrivals for hour $h$, day $d$, week $w$, and zone $z$, aggregate over the year.
@@ V Conclusion
In this short paper, we investigate the effectiveness of Tensor Chidori pseudoskeleton decomposition for anomaly detection in high-traffic areas of New York City. Specifically, we aim to capture temporal and spatial patterns in taxi arrival data. By focusing on central zones with significant activity, this method demonstrates the potential to capture sparsity and highlight urban regions with high demand.

### v3

# Robust Anomaly Detection via Tensor Pseudoskeleton Decomposition

@@ Abstract
Anomaly detection plays a critical role in modern data-driven applications, from identifying fraudulent transactions and safeguarding network infrastructure to monitoring sensor systems for irregular patterns. Traditional approaches—such as distance-, density-, or cluster-based methods, face significant challenges when applied to high-dimensional tensor data, where complex interdependencies across dimensions amplify noise and computational complexity. To address these limitations, this paper leverages Tensor pseudoskeleton decomposition within a tensor-robust principal component analysis framework to extract low-Tucker-rank structure while isolating sparse anomalies, ensuring robustness to anomaly detection. We establish theoretical analysis of convergence, and estimation error, demonstrating the stability and accuracy of the proposed approach. Numerical experiments on real-world spatiotemporal data from New York City taxi trip records validate the superiority of the proposed method in detecting anomalous urban events compared to existing benchmark methods. The results underscore the potential of Tensor pseudoskeleton decomposition to enhance anomaly detection for large-scale, high-dimensional data.
@@ I Introduction
To address these limitations, Tensor pseudoskeleton decomposition offers an alternative approach by selecting representative parts of the data, thereby preserving essential features while reducing redundancy. Tucker pseudoskeleton decomposition provides a structured decomposition that is both computationally efficient and robust [9, 10].
In this paper, we focus on anomaly detection within the tensor robust principal component analysis framework by leveraging a Tucker pseudoskeleton decomposition specifically tailored for high-dimensional datasets [9, 10]. By incorporating sparsity and regularization constraints, our method reduces sensitivity to anomalies, enabling more accurate and resilient detection of unusual patterns. The Tucker pseudoskeleton decomposition framework combines the strengths of Tucker decomposition’s structural insight with pseudoskeleton’s selective feature extraction while enhancing robustness against outliers [9].
@@ I-A Notations and definitions
In the realm of matrix algebra, the pseudoskeleton decomposition technique is a good alternative to SVD [12]. Specifically, this method entails selecting specific columns ${C}$ and rows ${R}$ from a matrix ${X}\in\mathbb{R}^{d_{1}\times d_{2}}$, and constructing a core matrix ${U}={X}(I,J)$. The matrix ${X}$ is then reconstructed through the product ${C}{U}^{\dagger}{R}$, under the condition that $\operatorname{rank}({U})=\operatorname{rank}({X})$. Expanding from matrices to tensors, the initial adaptations of pseudoskeleton decompositions applied a single-mode unfolding to 3-mode tensors [13]. To my best knowledge, the following are recent works on tensor pseudoskeleton decompositions or Tensor CUR Decompostions [9, 10, 14, 15, 16]. Furthermore, H. Cai, K. Hamm, and etc have presented rigorous theoretical results on the exact tensor pseudoskeleton decomposition [10, 14, 9]. For completeness, we present their work below.
Definition 4 (Tensor pseudoskeleton decompositions or Tensor CUR Decompostions [9, 10, 14]).
Consider a tensor $\mathcal{A}\in\mathbb{R}^{d_{1}\times\cdots\times d_{n}}$ with Tucker ranks $(r_{1},\dots,r_{n})$. Suppose that for each mode $i$, there exists a subset $I_{i}\subseteq[d_{i}]$ such that
$\mathcal{A}=\mathcal{R}\times_{i=1}^{n}\left(C_{i}U_{i}^{\dagger}\right),$
where $\mathcal{R}=[\mathcal{A}]_{I_{1},\dots,I_{n}},\quad C_{i}=[\mathcal{A}_{(i)}]_{:,J_{i}},\quad\text{and}\quad U_{i}=[\mathcal{C}_{(i)}]_{I_{i},:},\quad J_{i}=\bigotimes_{j\neq i}I_{j}.$
Theorem 1 [9, 10, 14]).
@@
For those interested in further details, It is recommended to read works [9, 10, 15, 16, 14, 17, 18].
@@ II Methodology
$\begin{split}\min_{\mathcal{R},{C}_{i},{U}_{i},\mathcal{S}}&\quad\|\mathcal{T}-\mathcal{L}-\mathcal{S}\|_{\mathrm{F}}\\ \text{subject to}&\quad\mathcal{L}={\mathcal{R}\times_{i=1}^{n}({C}_{i}{U}_{i}^{\dagger})}\\ &\quad\|\mathcal{S}\|_{\infty}\leq\alpha.\end{split}$
II-A TRPCA via Tensor Pseudoskeleton Decomposition
Algorithm 1 TRPCA via Tensor Pseudoskeleton Decomposition
1:Input: $\mathcal{T}\in\mathbb{R}^{d_{1}\times\cdots\times d_{n}}$: observed tensor; $(r_{1},\cdots,r_{n})$: estimated Tucker rank; $\varepsilon$: targeted precision; $\zeta^{(0)},\gamma$: thresholding parameters; $\{|I_{i}|\}_{i=1}^{n},\{|J_{i}|\}_{i=1}^{n}$: cardinalities for sample indices.
@@ II-A1 Step (I): Update Sparse Component $\mathcal{S}$
In this step, we update the sparse component $\mathcal{S}$ — which captures data outliers — using the technique described in [14, 19, 20]. Specifically, we apply an iterative decaying threshold within the hard thresholding operator $\mathrm{HT}_{\zeta}$ paired with $\gamma$, as described in [14, 21]. The hard thresholding operator $\mathrm{HT}_{\zeta}$ is defined as follows:
@@
This operator $\mathrm{HT}_{\zeta}$ effectively filters out entries with magnitudes less than or equal to $\zeta$, treating them as negligible. By applying this to the tensor $\mathcal{T}$, only values deemed significant (i.e., values exceeding the threshold) remain in the updated sparse component $\mathcal{S}$, thereby enhancing the sparsity of $\mathcal{S}$.
@@ II-A2 Step (II): Update Low-Tucker-rank Component $\mathcal{L}$
In this step, we aim to update the low-Tucker-rank component $\mathcal{L}$, which models the structured, low-rank part of the data tensor via tensor pseudoskeleton decomposition. The update process is divided into two key stages: subspace identification and projective reconstruction. To approximate the low-rank structure along each mode, we begin by extracting the mode-$i$ fibers from the residual tensor $\mathcal{T}-\mathcal{S}^{(k)}$, which represents the current estimate of the sparse component subtracted from the observed data tensor. The fibers are assembled into the matrix representation:
@@
This projection aligns the updated factor matrices along mode-$i$ with the estimated low-dimensional subspace. Using QR decomposition and projecting onto the selected subspace, the computational complexity for each mode is reduced from the cubic cost $\mathcal{O}(d_{i}^{3})$ to the more efficient: $\mathcal{O}(d_{i}r_{i}^{2}+r_{i}^{3}),$ where $d_{i}$ is the dimension along mode-$i$, and $r_{i}$ is the target Tucker rank. This reduction is particularly beneficial when the Tucker rank $r_{i}$ is significantly smaller than the mode dimension $d_{i}.$
@@ IV Numerical Experiments
In our experiments, we aggregate the data same as that in  [24] by counting the number of arrivals per zone over hourly intervals. To ensure statistical significance, we restrict our analysis to 81 central zones, which represent high-traffic areas and exclude zones with minimal activity. This selection reduces noise from sparsely populated zones and provides a more robust representation of NYC’s high-demand regions. With these parameters, we constructed a four-dimensional tensor $\mathbf{Y}$ with dimensions $24\times 7\times 53\times 81$. The modes of this tensor are defined as follows: the first mode corresponds to the 24 hours of a day; the second mode represents the 7 days of the week; the third mode encompasses the 53 weeks of the year; the fourth mode covers the 81 selected central zones in New York City. Thus, each entry in the tensor represents the count of taxi arrivals for hour $h$, day $d$, week $w$, and zone $z$, aggregate over the year.
@@ V Conclusion
In this short paper, we investigate the application of tensor pseudoskeleton decomposition for anomaly detection in high-traffic areas of New York City. Specifically, we aim to capture temporal and spatial patterns in taxi arrival data. By focusing on central zones with significant activity, the result demonstrates its possibility of tensor pseudoskeleton decomposition to remove sparsity and highlight urban regions with high demand.